Problem solving: to cook 4 cups of rice, you use 8 cups of water. To cook 10 cups of rice, you use 20 cups of water. Write an expression to show how many cups of water you should use if you want to cook c cups of rice.How many cups of water should you use to cook 5 cups of rice?
Answer by deborabr(181)

if you want cook 5 cups of rice you will need of 10 cups of water because the each cup of rice you use 2 cups of water
look
to 4 cups of rice are 8 cups of water and 8=4*2 or 8=4+4
is the same here 10 and 20 where 20=10*2 or 20=10+10
so we have 5 and 10 because 10=5*2 or 10=5+5
can express like this c+2w where c is the amount of cups of rice and w is the amount of cups of water